suppose a ball is dropped from a height of 30ft. and that each bounce reaches 3/8 of the previous height. how high will the ball boucnce on the second bounce?
Answer by venugopalramana(3286)

suppose a ball is dropped from a height of 30ft. and that each bounce reaches 3/8 of the previous height. how high will the ball boucnce on the second bounce?
FIRST BOUNCE HT.REACHED = 30*3/8
ON SECOND BOUNCE IT WILL REACH = 30*(3/8)(3/8)=270/64 =135/32 FT.